OpenStudy (nicole143):

Write the function that models each relationship. Find Z when X=4 and Y=8. Z varies jointly with X and Y. When X=2 and Y=2, Z=7. I have a table set up with the X's, Y's, and one Z but don't know where to go from there.. @mathmale

OpenStudy (mathmale):

"varies jointly" tells us that y is directly proportional to the values of both x and y. Symbolically, y=k*x*y, where k is just a constant of proportionality which we must find. We're told that z=7 when x=2 and y=2. substitute these three values into y=k*x*y, and then solve the resulting equation for k. Again, k is our constant of proportionality.

OpenStudy (nicole143):

Okay so I would be written as 2 = 7*2*2 ?

OpenStudy (nicole143):

*it

OpenStudy (mathmale):

If z= k * x * y, and z=7, x=2 and y=2, 7 = k * 2 * 2. Nicole, please solve this for k.

OpenStudy (nicole143):

Oh, I see. Sorry.

OpenStudy (mathmale):

So, z = ( ? )x*y

OpenStudy (nicole143):

k = 7/4 ?

OpenStudy (mathmale):

Yes, and so, if x= 4 and y = 8, z = ??

OpenStudy (nicole143):

16?

OpenStudy (nicole143):

Sorry, 56?

OpenStudy (btaylor):

56 is right.

OpenStudy (nicole143):

Okay, now that I have the equation used and the answer to Z, is that the problem solved? Or is there more?

OpenStudy (btaylor):

Yes, it is solved. You have Z=7/4(X)(Y), and found Z when x=2 and y=4. That is all that the question requires.

OpenStudy (nicole143):

Okay, thank you to both of you! Would you be willing to help me more?

OpenStudy (btaylor):

sure!

OpenStudy (nicole143):

Cool. Graph each equation. Identify the x- and y-intercepts and the asymptotes of the graph. y = 1/x I can put up pictures as we go for the graph.

OpenStudy (nicole143):

Do I just do a fill in table for x?

OpenStudy (btaylor):

You could.  Vertical asymptotes are whenever the denominator is 0 (here, x=0) Horizontal asymptotes are harder to describe...

OpenStudy (mathmale):

Nicole, BT is right on target, and I'm glad you two are working together. The function y=1/x is important and often seen in math. I'd suggest you memorize its shape, just as you would memorize the shape of the basic parabola